### v2

# Evaluating the Adversarial Robustness of Retrieval-Based In-Context Learning for Large Language Models

Abstract
With the emergence of large language models, such as LLaMA and OpenAI GPT-3, In-Context Learning (ICL) gained significant attention due to its effectiveness and efficiency. However, ICL is very sensitive to the choice, order, and verbaliser used to encode the demonstrations in the prompt. Retrieval-Augmented ICL methods try to address this problem by leveraging retrievers to extract semantically related examples as demonstrations. While this approach yields more accurate results, its robustness against various types of adversarial attacks, including perturbations on test samples, demonstrations, and retrieved data, remains under-explored. Our study reveals that retrieval-augmented models can enhance robustness against test sample attacks, outperforming vanilla ICL with a 4.87% reduction in Attack Success Rate (ASR); however, they exhibit overconfidence in the demonstrations, leading to a 2% increase in ASR for demonstration attacks. Adversarial training can help improve the robustness of ICL methods to adversarial attacks; however, such a training scheme can be too costly in the context of LLMs. As an alternative, we introduce an effective training-free adversarial defence method, DARD, which enriches the example pool with those attacked samples. We show that DARD yields improvements in performance and robustness, achieving a 15% reduction in ASR over the baselines. Code and data are released to encourage further research [^1] https://github.com/simonucl/adv-retreival-icl.
1 Introduction
Large Language models (LLMs) are revolutionising the field of NLP (Brown et al., 2020; Wei et al., 2022; Touvron et al., 2023). However, this also raises concerns about their robustness and trustworthiness. Although significant efforts are put into aligning LLMs for safety purposes, recent works still show LLMs can be vulnerable to adversarial inputs and jailbreaking (Zou et al., 2023; Xu et al., 2023b; Zeng et al., 2024). As a result, it is essential to understand LLM security properties to guide the direction of LLMs that are secure and robust. In this paper, we focus on examining a common LLM inference method, namely In-Context Learning (ICL) and its variants, on how they might be susceptible to perturbations and their robustness against adversarial attacks.
ICL is a common few-shot training method for prompting LLM for downstream tasks (Brown et al., 2020). Its variants include retrieval-based ICL models, which utilise retrievers to identify similar examples as demonstrations (Wang et al., 2023c; Li et al., 2023); or $k$NN-ICL, which compute label demonstrations using a nearest neighbours algorithm; have been adopted and shown promising performance improvements when compared to vanilla ICL (Shi et al., 2022; Xu et al., 2023a). However, the robustness of these methods against adversarial attacks received less attention and remains unclear. Prior work—which we review in Section 2—shows that LLMs can still be vulnerable to different types of input perturbations, such as task instructions (Zhu et al., 2023; Sun et al., 2023) or inputs (Wang et al., 2023a); however, the robustness of ICL-based methods to such perturbations is still under-explored.
We address (in Section 4) the following question: How sensitive are vanilla and retrieval-based ICL methods to perturbations to their test samples or demonstration instances, which is particularly important given that the primary motivation of ICL is to facilitate few-shot learning without training; should any methods exhibit excessive sensitivity, this would limit their applicability. To investigate this problem, we employ commonly used perturbation attacks—namely character-level, word-level, and sentence-level attacks. Furthermore, we propose new attack methods that target the demonstration instances and datastore examples. Our findings indicate that retrieval-based ICL methods show greater robustness against test-sample attacks; however, they can be more vulnerable when their demonstrations are adversarially perturbed. Detailed ablation studies further reveal that this vulnerability to adversarial attacks persists in newer models, such as Mistral (Jiang et al., 2023) and Gemma (Gemma Team et al., 2024). Many of these attacks can be transferable, both from the same family (LLaMA-2-7B $\rightarrow$ LLaMA-2-13B, 70B) or different family (LLaMA-2-7B $\rightarrow$ Mistral-7B). We showed Mixture of Expert (MoE) models still posed a similar adversarial threat to its dense version. This encourages further investigation on enhancing model robustness.
We further explore the advantages of utilising retrieval-based models to improve robustness (in Section 5), considering that existing defence strategies often rely on an adversarial training process (Li et al., 2021; Si et al., 2021). However, this process is memory intensive in the context of LLMs. Instead, we show that retrieval-based ICL methods, combined with adversarially augmented examples, yield more robust results than no augmentation methods while maintaining their predictive accuracy on in-distribution evaluation sets. Therefore, we propose DARD, a training-free adversarial defence method which leverages adversarially perturbed samples to augment the retrieval pools, potentially improving the models’ robustness. The paper is summarized in the Figure 1.
Our main contributions are summarized as follows.
  • We first perform a comprehensive and in-depth evaluation on variants of ICL: vanilla, $k$NN-ICL, and retrieval-based ICL methods against Test-Sample, Demonstration and Datastore Adversarial Attacks.
  • We show that Retrieval-based models can bring positive results on robustness when test samples are under adversarial attack, but it further reduces the robustness while demonstrations are perturbed. We show these attacks persist on larger models and are transferable.
  • We propose DARD, a novel training-free adversarial defence method which leverages adversarially perturbed examples to improve the robustness.
2 Related Work
Adversarial attack and defences
Adversarial attack has always been a topic in deep neural networks, where previous works showed neural networks can be vulnerable to adversarial examples that are crafted by adding imperceptible but adversarial noise on natural examples (Szegedy et al., 2013; Evtimov et al., 2017). In the real world, adversarial robustness is important for the application of a method (Tramèr et al., 2017).
Several works studied this problem and the main finding is that LMs struggle to answer correctly and factually under attacks and propose their defence via label smoothing or prompt tuning (Yang et al., 2022; Raman et al., 2023). It is important to note, however, that these experiments were mainly conducted using relatively small LMs such as BERT, RoBERTa. Recently, several works demonstrated LLM can still be vulnerable to these type of attacks, by showing the vulnerability of attacking the task instructions (PromptBench (Zhu et al., 2023)), or perform adversarial attacks on zero-shot settings and investigate the transferability of attacks from open-sourced models (i.e. Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023)) to close-sourced (i.e. GPT-3.5 and GPT-4) (DecodingTrust (Wang et al., 2023a)). However, previous study (Gu et al., 2022) has shown that the LLMs are less sensitive to prompt / instruction variation when few-shot examples are provided in context (a.k.a ICL). However, there is still a lack of systematic analysis on this case which is the main contribution of this paper. Wang et al. (2023b) shows that attacking demonstrations can still be an effective way for adversarial attacks where LLMs are still vulnerable. Therefore, we aim to extend and evaluate the adversarial attack and defence performance on retrieval-based models.
ICL Sensitivity
The existing literature on ICL analysis can be broadly divided into two streams, each focusing on different aspects. The first stream explores the influencing factors of ICL based on input perturbation, such as the order (Min et al., 2022), the formatting (Yoo et al., 2022; Lu et al., 2022) and the selection of the demonstration (Liu et al., 2022). Designing proper demonstration construction strategies (Ye et al., 2023; Li et al., 2023) could bring clear boosts to the ICL performance, while Chen et al. (2023) analyses the sensitivity of ICL via perturb on instructions and example ordering, and our work is a complement of their work on sensitivity.
This paper provides a different view of when different ICL methods fail and when retrieval-based ICL can be a better approach for robustness.
Retrieval-based LLM
Retrieval-based LLM has shown great success in lifting LLM performance and faithfulness and reducing hallucination, where the idea is to gather related articles or demonstrations from a collection of data. Specifically, we are interested in whether they can be used to improve the robustness against perturbations (Lin & Lee, 2024). Meanwhile, many papers also focus on analysing the robustness of RAG methods on irrelevant contexts (Shi et al., 2023; Asai et al., 2024), in which we adopt similar methods in our context for analysing how retrieval-based ICL being sensitive to irrelevant contexts.
Figure 1: Overview of the paper. We visualize our seven adversarial attacks in (a), (c) and (d) (only 3 shots are used in the plot for display purposes). And our adversarial defence method, DARD, is showcased in the top right corner (Plot (b)).
3 Problem Formulation
3.1 Background
In this work, we consider several types of ICL methods, namely Vanilla ICL, and two types of retrieval-based ICL: $k$NN-ICL (Xu et al., 2023a) and Retrieval-based ICL (Wang et al., 2023c; Li et al., 2023). Retrieval-based ICL methods demonstrated promising performance improvements over vanilla ICL ones; however, their robustness against adversarial perturbations remains unexplored.
Vanilla ICL
is the most commonly used few-shot learning method in LLM. Given a training data set $\mathcal{T}=\{(x_{i},y_{i})\}$, where $x_{i}\in X$ denote the instances and $y_{i}\in Y$ represent their labels, the concatenation of the $k$-shot input-label pairs is used to generate the prompt $P$ with respect to a test instance $x_{\text{test}}$:
$P=\mathcal{I}\oplus\pi(x_{1},y_{1})\oplus\pi(x_{2},y_{2})\oplus\dots\oplus\pi(x_{k},y_{k})\oplus\pi(x_{\text{test}},*)$ (1)
where $\mathcal{I}$ denotes the task instructions, $\pi$ denotes the templates and verbaliser to turn inputs into prompts (details for each dataset are in Appendix A.2), $\oplus$ is a concatenation operator, and $|Y|$ is the number of distinct labels. The label prediction $y_{\text{test}}$ for the test instance $x_{\text{test}}$ is then computed as:
$y_{\text{test}}=\arg\max_{y\in Y}p(y|P;\theta_{\text{LLM}})$ (2)
$k$NN-Prompting
was first introduced by Shi et al. (2022) to perform zero-shot inference and extended in later work by Xu et al. (2023a) to support few-shot inferences, simulating the ICL. Throughout the paper, we follow the implementation of Xu et al. (2023a) and call it $k$NN-ICL. The idea is to map each training instance into a latent representation using a LLM and cache the last token distribution into a datastore. At inference time, the prediction is based on the interpolation between the LLM prediction and the $k$-nearest neighbours label distribution in the datastore. Specifically, for each training instance $x_{i}$, Xu et al. (2023a) concatenate it into prompt $P_{i}$ together with $k$ in-context examples, where $k=|Y|$:
$P_{i}=\pi(x_{1},y_{1})\oplus\dots\pi(x_{|Y|},y_{|Y|})\oplus\pi(x_{i},*).$ (3)
By querying the LLM using $P_{i}$, we obtain a distribution over tokens $p(v|P_{i},\theta)$. Rather than mapping it back to label space $Y$, the distribution is cached as the key representation for the instance $x_{i}$:
$\mathbf{k_{i}}=p(v|P_{i};\theta_{\text{LLM}}),$ (4)
which is mapped to the label $y_{i}$. The entire datastore thus consists of $\{(\mathbf{k_{i}},y_{i})\}$ pairs, where $K=\{\mathbf{k_{i}}\}_{i}$ denotes the set of keys. At inference time, we construct the same prompt and obtain the distribution $P_{\text{test}}=p(v|P_{\text{test}};\theta_{\text{LLM}})$. We then match the distribution against cached $K$ in the datastore, where standard $KL$-divergence is used to measure the distance:
$D_{\text{KL}}(P_{\text{test}}||k_{i})=\sum_{v}p(v|P_{\text{test}};\theta_{\text{LLM}})\log\frac{p(v|P_{\text{test}};\theta_{\text{LLM}})}{p(v|P_{i};\theta_{\text{LLM}})}$ (5)
The predictions are then calculated by interpolating the LLM output and its $m$ nearest neighbours distribution, where $\text{NN}_{m}(*,K)$ denotes the set of $m$ nearest neighbours in $K$:
$\hat{y}_{\text{pred}}=\arg\max_{y\in Y}\left[(1-\alpha)\cdot p(y|P_{\text{test}};\theta_{\text{LLM}})+\alpha\cdot p\sum_{i\in\text{NN}_{m}(P_{\text{test}},K)}\mathbb{I}(y_{i}^{a}=y)\right].$ (6)
Following Xu et al. (2023a), in this work, we use $\alpha=0.2$ and $m=k/2$.
R-ICL
Retrieval-augmented ICL (Lin et al., 2022) was proposed to mitigate the sensitivity of ICL methods to the choice and order of the in-context examples: rather than randomly sampling $(x_{i},y_{i})$, R-ICL use a retriever to identify the most similar examples to the test instance $x_{\text{test}}$ from the training set $\mathcal{T}$. In this paper, we follow Li et al. (2023) and consider the following retrievers: BM25 (Robertson & Zaragoza, 2009), SBERT (Reimers & Gurevych, 2019) and Instructor (Su et al., 2022); more details are available in Appendix A.1.
3.2 Attack Methods
This paper aims to study how perturbations can affect the performance of different ICL methods through adversarial attacks. We classify our 7 attacks into 3 categories: Test Sample Attack, Demonstration Attack, and DataStore Attack.
Test Sample Attacks
We focus on the three different types of attack: typo-based perturbations — TextBugger (TB) (Li et al., 2018); embedding-similarity-based perturbations — TextFooler (TF) (Jin et al., 2019); and context-aware perturbations — BERT-Attack (BA) (Li et al., 2020).
Demonstration Attacks
For demonstration attacks, we refer to adversarial attacks where the training demonstrations are adversarially perturbed. We consider this type of attack to investigate the sensitivity of LLMs to adversarial perturbations to the training demonstrations and their influence on the generalisation properties of the models.
Datastore Attacks
One of our objectives is to analyse the sensitivity of retrieval-based ICL methods to adversarial perturbations. Therefore, we consider Irrelevant Context Attacks (Irr.), which work by contaminating the demonstration pools with irrelevant contexts. Our Irrelevant Context Attacks extended Min et al. (2022) — we replace 50% of the original demonstrations in the training data with Out-of-Distrubtion (OoD) examples. Further details and analyses are available in Appendix A.3.
4 Experiments
4.1 Settings
Datasets
To assess the robustness, we selected six sentiment and multiple-choice text classification datasets: SST2 (Socher et al., 2013), RTE (Dagan et al., 2005), CR (Hu & Liu, 2004), MR (Pang & Lee, 2005), MNLI-matched (Williams et al., 2018), and TREC (Voorhees & Tice, 2000). We provide further details, including prompt templates and dataset statistics, in Appendix A.2. Unless explicitly mentioned, LLaMA-2-7B was mainly used as the base model for our experiments (Touvron et al., 2023). For experiments involving randomness (i.e., Vanilla ICL and $k$NN-ICL), we conducted trials with three different random seeds to ensure the reliability of our findings.
Evaluation
To evaluate and compare the vulnerabilities across different types of attacks, in line with Wang et al. (2023a); Zhu et al. (2023), we assess the model’s performance in a benign (attack-free) context (Clean) and calculate the Attack Success Rate (ASR). The ASR is defined as Clean AccuracyAttack AccuracyClean $\frac{\text{Clean Accuracy}-\text{Attack Accuracy}}{\text{Clean Accuracy}}$, where Attack Accuracy denotes the model’s accuracy under attack.
4.2 Results
The 8-shot attack results are presented in Table 1 with 8-shot attacks, and detailed results are available in Appendix B.1.
Overview:
Retrieval-based models exhibit higher Clean accuracy: 3.27% higher than vanilla ICL with the best performed SBERT retriever; similar finding with Rubin et al. (2021) that related demonstrations enhance the performance. For adversarial attacks, our proposed adversarial attack methods, Swap-Labels and Swap-Labels (Fix), are effective techniques for manipulating LLM predictions. On the other hand, irrelevant context attacks have a negligible effect on performance. We summarise two major findings about the adversarial robustness across different ICL methods.
Tasks Method Clean TB TF BA AdvICL S-L S-L (Fix) Irr. Avg
SST-2 ICL 92.66 37.29 43.69 68.40 33.64 92.33 62.99 1.69 48.58
$k$NN-ICL 92.01 41.01 45.45 61.66 79.05 67.28 66.00 0.0 51.49
$\text{R}_{\text{BM25}}$-ICL 94.61 43.39 45.94 57.33 25.63 99.88 63.88 0.85 46.7
$\text{R}_{\text{SBERT}}$-ICL 95.18 33.01 33.01 56.26 27.21 99.64 66.84 6.09 46.01
$\text{R}_{\text{Instructor}}$-ICL 94.84 38.33 32.62 58.65 28.22 99.54 69.32 6.73 47.63
RTE ICL 73.04 82.54 75.62 97.86 14.32 91.59 58.31 11.04 61.61
$k$NN-ICL 70.52 90.61 80.37 95.43 96.13 63.66 63.15 6.83 70.88
$\text{R}_{\text{BM25}}$-ICL 71.48 68.69 63.14 90.4 18.77 92.42 60.60 15.15 58.45
$\text{R}_{\text{SBERT}}$-ICL 72.20 83.50 83.01 98.12 27.49 95.50 64.50 15.50 66.80
$\text{R}_{\text{Instructor}}$-ICL 73.29 78.33 72.41 96.67 28.08 90.64 66.00 10.36 63.21
MR ICL 92.67 36.41 46.72 67.52 22.86 94.49 75.43 0.15 49.08
$k$NN-ICL 91.77 39.60 47.18 65.95 78.99 71.81 72.90 2.78 54.17
$\text{R}_{\text{BM25}}$-ICL 92.50 37.95 45.84 65.95 25.10 99.68 63.89 1.51 48.56
$\text{R}_{\text{SBERT}}$-ICL 92.40 38.53 46.10 64.52 28.83 99.78 60.06 0.0 48.26
$\text{R}_{\text{Instructor}}$-ICL 92.6 37.47 47.30 66.22 26.26 98.06 65.33 0.43 48.72
CR ICL 91.31 35.54 52.43 71.16 22.8 99.90 91.87 1.22 51.56
$k$NN-ICL 91.87 19.87 39.13 58.78 74.07 67.33 57.74 6.03 44.13
$\text{R}_{\text{BM25}}$-ICL 93.09 29.14 47.43 65.72 33.20 100.00 56.29 7.30 48.44
$\text{R}_{\text{SBERT}}$-ICL 93.62 31.25 46.30 71.59 33.81 100.00 62.32 4.76 50.00
$\text{R}_{\text{Instructor}}$-ICL 93.88 34.55 39.38 53.82 35.81 100.00 64.88 4.02 47.49
MNLI-mm ICL 53.63 68.49 57.92 40.85 32.48 88.81 43.84 1.73 47.73
$k$NN-ICL 55.89 65.61 52.10 51.30 51.30 62.39 56.00 10.14 49.83
$\text{R}_{\text{BM25}}$-ICL 57.3 68.41 59.16 40.98 38.45 95.99 63.89 0.35 52.46
$\text{R}_{\text{SBERT}}$-ICL 57.8 69.03 55.07 43.27 37.01 99.31 63.55 2.77 52.86
$\text{R}_{\text{Instructor}}$-ICL 54.30 66.48 51.93 36.06 37.83 100.00 69.67 2.58 52.08
TREC ICL 76.07 61.09 57.50 60.65 18.69 65.56 40.58 8.19 44.61
$k$NN-ICL 77.87 59.68 54.54 50.02 83.65 58.74 52.61 10.88 52.87
$\text{R}_{\text{BM25}}$-ICL 79.20 50.00 50.76 57.07 28.91 55.56 9.60 3.79 36.53
$\text{R}_{\text{SBERT}}$-ICL 87.8 43.74 45.56 47.15 34.61 53.08 10.02 9.11 34.75
$\text{R}_{\text{Instructor}}$-ICL 86.5 45.43 70.87 47.51 31.31 44.97 3.35 7.28 35.82
Avg ICL 79.90 53.56 55.65 67.74 24.13 88.78 59.84 4.00 50.53
$k$NN-ICL 79.98 52.73 53.13 63.86 77.20 65.2 61.40 6.11 54.18
$\text{R}_{\text{BM25}}$-ICL 81.36 49.6 52.04 62.91 27.68 90.59 53.02 4.82 48.52
$\text{R}_{\text{SBERT}}$-ICL 83.17 49.84 51.51 63.48 31.49 91.22 54.55 6.37 49.78
$\text{R}_{\text{Instructor}}$-ICL 82.57 50.10 52.42 59.82 31.25 88.87 56.42 5.23 49.16
Table 1: Adversarial Attack results with 8-shot ICL. Clean refers to the Benign accuracy, while all other columns refer to the Attack Success Rate (%) under corresponding Adversarial Attacks. The last column Avg refers to the mean ASR across 7 attacks. The higher the ASR, the lower the robustness of models against the attack.
Figure 2: Attack Success Rate (ASR) across shots among different ICL methods. We aggregated the results for ICL and $k$NN-ICL across 3 seeds and R-ICL results across 3 retrievers.
Finding 1: Retrieval-based models are more robust against test-sample attacks; however, it could fall short under demonstration attack.
[…]
On the other side, we see an increase in ASR for Retrieval-based ICL when demonstrations are under attack, demonstrating a 1%-6% increase in ASR compared with vanilla ICL. This demonstrates the deficit of retrieval-based models that can sometimes overly rely on the retrieved demonstrations and are more vulnerable when demonstrations are perturbed.
Finding 2:
$k$NN-ICL is highly sensitive to adversarial perturbations on test samples and demonstration.
In Table 1, $k$NN-ICL shows the highest ASR in Demonstration Attack, specifically showing the highest vulnerability to AdvICL, 40%-50% higher than ICL and R-ICL. Recall that for $k$NN-ICL, demonstrations and test samples are mapped into the hidden space to compute the label distribution with nearest neighbours. Any perturbations on the demonstration might notably affect their position in the hidden space and alter the label distribution significantly. Also, in Figure 2, $k$NN-ICL exhibits a consistently ascending trend in ASR as the number of examples increases, with 15% higher than the other methods. This further illustrates that incorporating nearest neighbours algorithms inevitably includes more noise, thereby increasing vulnerability.
Overall, our experiment results show that few-shot methods in LLM could still be vulnerable to adversarial perturbations, leading to a 50% drop in performance on average. Pairing with previous works on jailbreaking leads to the need for a continuous effort to improve the robustness of LLMs. We include quantitative results in the Appendix B.2.
4.3 Ablation Study
Figure 3: Attack Success Rate (%) for adversarial attacks across various models, based on experiments conducted on the RTE dataset with 8-shot demonstrations. The results are based on R-ICL and involve the aggregation of results from retrievers: BM25, SBERT, and Instructor.
In this section, we extend our experiments to four other models: LLaMA-2-13B, larger version of 7B model; Vicuna-7B (Chiang et al., 2023), instruction-tuned from LLaMA-2-7B; Mistral-7B (Jiang et al., 2023) and Gemma-7B (Gemma Team et al., 2024), aiming to provide a view for models from different families to prevent potential bias posed within the pretraining data for LLaMA models. These results allowing us to analyze the vulnerability of modern models with better coverage. We conducted the experiments on RTE datasets with 8-shot demonstrations. Even larger models such as LLaMA-2-70B or Mixtral-8$\times$7B (Jiang et al., 2024) are ignored due to attacking them is expensive and slow. Instead, we leave them in the following section for attack transferability (§4.4).
The results are shown in Figure 3 for R-ICL (corresponding results for ICL are presented in Figure 5). It can be observed that most adversarial attacks remain effective across all models, including the Swap-Labels attack we proposed. Thus, it still posed a major threat on modern LLMs. Specifically, Vicuna-7B tends to be more sensitive when attack on text sample (TextBugger) or when demonstrations are under attack (AdvICL); while Gemma-7B—most recently released model—shows a higher vulnerability when the context or label in demonstrations is adversarial perturbed, compared to other base models. Notably, the irrelevant context attacks become less vulnerable to more recent release models, especially Vicuna, which shows nearly no effect after being tuned with instructions, showing improved capability in the latest LLMs for filtering OoD contexts (Shi et al., 2023).
4.4 Attack Transferability
Figure 4: Analysis of the attack transferability for Retrieval-ICL on larger variant models within the same family: (Left) the LLaMA family; (Mid) the Mistral family; and (Right) across models from different families. The models’ orders are sorted by their parameter sizes and release date. The models highlighted in bold are the models being attacked.
In this section, we would like to answer two questions: Are these attacks transferable across different models, and can larger models mitigate the vulnerability to these attacks? To this end, we conducted two sets of experiments: (1) transfer attacks within the same model family from small models to their large variants, by testing on LLaMA-2 and Mistral famililes; (2) whether attacks still be effective across models from different family, by transferring attacks to Vicuna-7B, Mistral-7B and Gemma-7B, following the same settings as Section 4.3. Our experiments focused on the RTE dataset, chosen for its pronounced drop in robustness when subjected to these attacks. Due to memory constraints, both Mixtral-8$\times$7B and LLaMA-2-70B were run with 8-bit quantization (Frantar et al., 2022), while the remaining models were in full precision. Hong et al. (2024) tested on AdvGLUE++ and showed that quantized LLMs have negligible drops in adversarial robustness to their source model.
The results of the transfer attack with R-ICL are shown in Figure 4 (corresponding results for ICL are presented in Figure 6). Most attacks are transferable and achieve an ASR of 15%-40% on similar scale models; except for AdvICL, which cannot transfer between models. Also, Swap-Labels attacks are shown to be less effective when targeting instruction-tuned models (e.g. Vicuna and Mistral-7B-Instruct), pointing to the instruction-tuning process leading models to be less sensitive to its demonstration (Sun et al., 2023). Specifically, the results shown in Figure 4(L) confirm that larger models tend to be more robust, as indicated by a reduction of 6% on average in ASR between the 13B and 70B models. However, the MoE variants of Mistral-7B-Instruct exhibit a similar or even higher ASR compared to its dense version, as illustrated in Figure 4(M). This finding deviates from previous work by Puigcerver et al. (2022), which investigates adversarial robustness in MoE, showing Vision MoE models are less vulnerable to adversarial attacks. We believe this discrepancy might be subject to the difference between on how MoE is applied in NLP (token) and Vision (image-patch) models. Remember that in the training of MoE models, experts are specialised to handle different tasks (Xue et al., 2024); perturbations in the input texts might change the router output and be routed to experts that are less capable for the targeted tasks, leading to more vulnerability and changes in the output. We hope our findings encourage future work to study the robustness of MoE models, both from the training and router perspectives, given their widespread use nowadays.
5 Defence Methods: DARD
In this section, we propose an easy yet effective method to enhance the robustness of retrieval-based models against adversarial attacks on test samples. Specifically, we introduce an adversarial defence method named Demonstration Augmentation Retrieval Defences (DARD). This approach performs adversarial augmentation on the training data and mixes them into the retrieval bases for R-ICL. The initiative is that adversarial training in LLM can be costly in memory, and Xie et al. (2021); Dai et al. (2022) showed that in-context learning can be seen as an implicit finetuning method.
5.1 Methodology
The methods begin by adversarially perturbing the training examples, similar to previous approaches. Our experiments use the RTE dataset, focusing on an 8-shot setting for this section. For each of the 872 test samples in RTE, the retriever selects the 8-shot most similar examples for demonstrations. After deduplication, we are left with 7,293 distinct examples. These are concatenated with one-shot examples, and adversarial attacks (TF, TB, and BA) are performed on the extracted training samples. Only the successful examples are retained, resulting in 16,206 perturbed examples. In the adversarial training paradigm, such perturbed examples would typically be used to fine-tune models. However, our approach, DARD, reintegrates these examples into the training demonstration for retrieval during the inference stage. It is important to note that we impose a constraint whereby each example and its perturbed examples can be retrieved no more than once in same prompt for ICL. This is because we observed that most retrieved samples were variants of the same examples with perturbations without this constraint, which reduced both performance and robustness.
Defence Shots Clean TB TF BA Avg
$\hookrightarrow$ No Defence 8 71.48 68.69 63.14 90.4 74.08
Augmentation
$\hookrightarrow$ Random Addition 8 72.01 67.75 67.76 91.6 75.70
$\hookrightarrow$ Random Deletion 8 69.18 65.72 62.78 85.96 71.49
DARD
$\hookrightarrow$ R-ICL (BM25) 8 74.39 58.30 51.43 79.32 63.01
$\hookrightarrow$ R-ICL (SBERT) 8 72.16 58.53 44.46 72.60 58.53
$\hookrightarrow$ R-ICL (Instructor) 8 71.26 69.00 61.81 84.31 71.70
Adversarial Training 8 77.22 62.22 41.96 77.22 60.47
Table 2: DARD for adversarial defences. We show the Clean Accuracy and the ASR. For Random Addition and Deletion baselines, BM25 is the retriever as it performs the best.
5.2 Results
For baselines, we compared with No Defence, Random Augmentation (Addition, Deletion) inspired by Robey et al. (2023), which are smoothing methods used to defend against attacks. We also compared with adversarially trained models following vanilla ICL setup. We show the results in Table 2 (detailed results are in Table 18). Although the adversarially trained models show a higher clean accuracy as it’s directly fine-tuned on RTE data, our methods with SBERT outperform it by  2% reduction in ASR; both BM25 and SBERT variants of DARD have significant improvement on robustness against No Defence baseline.
6 Summary
While ICL sensitivity is a widely known issue, our paper provides an alternative view of the problem through adversarial attacks. In summary, our paper conducts a comprehensive analysis of the effects of perturbations on modern few-shot methods for LLMs inference and also proposes new adversarial attacks targeted at ICL models. Besides their performance, the findings show the other side regarding retrieval-based models related to their robustness. Furthermore, we also demonstrate the potential of retrieval-based models in data augmentation strategies for adversarial defense.
There are many open questions for future work.One of the most profound findings is that our study on attack transferability demonstrates that larger models generally exhibit better robustness; however, this does not necessarily hold true for Mixture of Experts (MoE) models. Given the recent surge in their popularity, it is crucial to investigate the conditions under which MoE models may fail and to understand the reasons behind their vulnerabilities.
Appendix A Experiment Settings
This section discusses the details of the ICL methods used, including their hyperparameters, and also the verbalizer and prompt templates for classification tasks. The details on each attack are also provided for reproducibility. Code are included as supplementary materials and will be publicly available upon the paper is accepted.
A.1 Details about the ICL methods
$k$NN-ICL the exact formulation follows the original paper (Xu et al., 2023a).
R-ICL We give more details about the retriever used throughout the paper here:
  1. 1.
    BM25 (Robertson & Zaragoza, 2009): A prevailing sparse retriever. For BM25, we follow Luo et al. (2023) and use uncased BERT wordpiece tokenization (Devlin et al., 2019).
  2. 2.
    SBERT (Reimers & Gurevych, 2019): We use the Sentence-BERT as the dense demonstration retriever. Specifically, we used all-MiniLM-L6-v2 to encode the test input and training set’s inputs, and retrieve the examples with the most similar input as demonstrations.
  3. 3.
    Instructor (Su et al., 2022): Instructor is a recently proposed competitive text embedding model trained on 330 tasks with instructions. By providing specialized instruction, it can serve for demonstration retrieval. We conduct experiments on its released large-size model 333hkunlp/instructor-large.
A.2 Templates and verbaliser for datasets
See Table 3 for the used templates and verbaliser for mapping discrete labels into label space (Adopted from Min et al. (2022) and Lu et al. (2022)).
Task Template Verbaliser
SST2 Review: contains no wit , only labored gags 0: negative,
Sentiment: negative 1: positive
Review: the film is powerful , accessible and funny .
Sentiment:
RTE A man is due in court later charged with the murder 26 years ago of a teenager whose case was the first to be featured on BBC One’s Crimewatch. Colette Aram, 16, was walking to her boyfriend’s house in Keyworth, Nottinghamshire, on 30 October 1983 when she disappeared. Her body was later found in a field close to her home. Paul Stewart Hutchinson, 50, has been charged with murder and is due before Nottingham magistrates later.” 0: false, 1: true
The question is Paul Stewart Hutchinson is accused of having stabbed a girl. True or False?
The Answer is: false
For women earning 22,000 a year, the total pay accumulated after six months maternity leave would be just 5,300 in the UK and 5,850 in Ireland. Entitlements in Germany would also be relatively low, at 5,900, along with those in France, Spain and the Netherlands, all at 6,750. At the other end of the scale, pay received after six months leave in Italy would be 9,150 while in Denmark and Norway it would be as much as 11,000.
The question is Maternity leave varies in Europe. True or False?
The Answer is:
MR Review: ”you might say tykwer has done all that heaven allows , if you wanted to make as anti-kieslowski a pun as possible . suffice to say its total promise is left slightly unfulfilled .” 0: negative, 1: positive
Sentiment: negative
Review: an alternately raucous and sappy ethnic sitcom . . . you’d be wise to send your regrets .
Sentiment:
CR Review: it ’s not as stylized as a sony or samsung . 0: negative,
Sentiment: negative 1: positive
Review: i went out and got the canon today .
Sentiment:
MNLI Instruction: Please identify whether the premise entails the hypothesis. The answer should be exactly ’yes’, ’no’ or ’maybe’. 0: yes, 1: maybe, 2: no
Premise: We serve a classic Tuscan meal that includes a Florentine terrine made with dick and chicken livers .
Hypothesis: We serve a meal of Florentine terrine .
Prediction: yes
Premise: After a lifetime of trials, Donna not only earned her GED at Goodwill, she earned a job here.
Hypothesis: Donna went through Goodwill to get her GED, but was still unemployed.
Prediction:
TREC Question: How did serfdom develop in and then leave Russia ? 0: expression,
Type: description 1: entity,
2: description,
Question: What is Shakespeare ’s nickname? 3: human,
Type: 4: location,
5: number,
Table 3: Templates and verbaliser used across the experiments. These are minimum cases with only one demonstration example for illustration.
A.3 Attack Details
We implemented our proposed attack methods and conducted all our experiments using the TextAttack framework (Morris et al., 2020), which is model-agnostic and can be easily tested on any model. Here are the details regarding to each attack method:
  1. 1.
    TextBugger: We employ the TextBugger (Li et al., 2018), which determines the importance of words and replaces key words to conduct the attack. We use GloVe embeddings to retrieve nearby words of the words to be replaced for the attack.
  2. 2.
    TextFooler: We also evaluated another popular word-level method, TextFooler (Jin et al., 2020), which is similar principle to TextBugger. However, TextFooler greedily searches for a large number of nearby words in the embedding space for each word, as long as they meet some constraints on embedding similarity and sentence quality. Additional constraints require the replacement words to match the Part-Of-Speech (POS) of the original words.
  3. 3.
    BERT-Attack: We used the popular character-level attack, BERT-Attack (Li et al., 2020), which utilizes BERT to generate candidate replacement words based on the importance of words. Then, by calculating the impact of each replacement word on the classification outcome, the replacement word that has the most significant impact on the classification result is chosen as the final replacement.
  4. 4.
    AdvICL: We mostly follow the implementations by Wang et al. (2023b), except for limiting the attack percentage to be 15%. Also, for RTE and MNLI datasets, which consist of both premise and hypothesis, we only allow perturbations on the hypothesis as it’s closer to the realistic cases, and the premise is usually much longer than the hypothesis.
  5. 5.
    Swap-Labels/Swap-Labels (Fix): For both of the attacks, we reused the GreedyWordSwapWIR implementation in textattack framework. The attacks start by measuring the importance of each label in the demonstration. It is done by first replacing each of the labels with a meaningless word, fetching it into the LLM to get the label distribution, and computing the difference to the original label distribution. Based on this, the Swap-Labels attack focuses on the most important labels first, $y_{i}$, by replacing it with other labels from the label set, namely $y_{i}^{\prime}\in\{Y/y_{i}\}$. If that reduces the LLM output probability on the correct label, the attack models continue to attack the subsequent most important labels with the perturbed examples with $y_{i}$ replaced by $y_{i}^{\prime}$; otherwise, it keeps $y_{i}$ and attack another labels. This process iterates until the attack is successful or the maximum perturbations allowed are reached.
    The only difference between Swap-Labels (Fix) and Swap-Labels is that an additional constraint is imposed: the label distribution on the attacked samples should be the same as the original distribution.
  6. 6.
    Irrelevant Context Following Min et al. (2022), CC-news (Hamborg et al., 2017) is used as the corpus providing OoD sentences. Sentence length is considered to select the most relevant sentences to the original demonstration. The contamination rate is 50%.
Examples of each attack are demonstrated in Table 5-11. Code will be released to facilitate further research in this field.
A.4 Reproducability
We conducted all attack experiments using either two A100-40GB GPUs or one A100-80GB GPU. The batch size is 8. For ICL and $k$NN-ICL that consists of randomness, we conducted experiments with seed $\{1,13,42\}$. Most attacks took between 1 and 2.5 hours to run. The exception was the 16-shot demonstration attack, which usually lasted 5 hours or longer. We used Flash-Attention 2 (Dao, 2023) to speed up the experiments, which gave us a 2 to 2.5 times speed increase. The speedup significantly helped our comprehensive experiments.
Appendix B Complete Results
Figure 5: Attack Success Rate (%) for adversarial attacks across various models, based on experiments conducted on the RTE dataset with 8-shot demonstrations. The results are based on ICL and involve aggregating results from 3 shots.
Figure 6: Transferable attack for ICL on larger variant models within the same family: (L) the LLaMA family; (M) the Mistral family; and (R) across models from different families. The models’ orders are sorted by their parameter sizes and release date. The models highlighted in bold are the adversarial models under attack.
B.1 Main results
We include the complete results breakdown in Table 12-17. The results shown in those tables are the Benign accuracy and Attack accuracy.
B.2 Quantitative Results
We conducted a human study to evaluate the feasibility of adversarial attacks and estimate human performance against them. Specifically, we provided 100 input examples to students and asked them to determine the correct label for each given sentence (answer options: true or false). We used the RTE dataset and instructed participants to give their answers solely based on the provided context. We assessed various types of attacks.
In this set of 100 sentences, the baseline model we used, $\text{R}_{\text{Instructor}}$-ICL , achieved an accuracy of 88% on the original, unaltered sentences. First, we observed that word-level attacks on test examples had a minor impact on human judgment, with an average decrease of only 3% (Table 4). Attacks such as advicl and irrelevant attacks resulted in decreases of only 7% and 0%, respectively. However, these attacks could significantly affect retrieval models, causing attack successful accuracies (ASR) ranging from 15% to 95%, indicating that the models are more fragile and sensitive than anticipated.
For swap-label and swap-label (fix) attacks, human performance decreased by 25% and 13%, respectively. This suggests that attacks manipulating the label of demonstrations indeed caused confusion for humans, demonstrating their effectiveness.
TB TF BA AdvICL S-L S-L (Fix) Irr.
Perturbed Examples 84% 87% 82% 81% 63% 75% 88%
Table 4: Summary of the responses from the human evaluation. The total number of evaluated examples is 100. The accuracy under unperturbed conditions on these 100 examples is 88%.
Unperturbed only a few mag-lev trains have been used commercially such as at the birmingham airport in the uk.,
The question is: maglev is commercially used. True or False?
The answer is: ture.
they, too, have not produced a practical, commercially acceptable maglev.
The question is: maglev is commercially used. True or False?
The answer is: false
two weeks ago, china became the first nation to operate a maglev railway commercially, when officials inaugurated a 30-kilometer-long line between downtown shanghai and the city’s airport.
The question is: maglev is commercially used., True or False?
The answer is: true.
the m-2000 uses a commercially manufactured nbti superconductor, similar to that used in the m-2000 maglev magnets.,
The question is: maglev is commercially used. True or False?
The answer is: false
it appears that the super-conducting maglev system is technically ready to be used commercially as a very high-speed, large-capacity transportation system.
The question is: maglev is commercially used. True or False?
The answer is false.
Perturbed only a few mag-lev trains have been used commercially such as at the birmingham airport in the uk.,
The question is: maglev is commercially used. True or False?
The answer is: ture.
they, too, have not produced a practical, commercially acceptable maglev.,
The question is: maglev is commercially used. True or False?
The answer is: false
two weeks ago, china became the first nation to operate a maglev railway commercially, when officials inaugurated a 30-kilometer-long line between downtown shanghai and the city’s airport.
The question is: maglev is commercially used., True or False?
The answer is: true.
the m-2000 uses a commercially manufactured nbti superconductor, similar to that used in the m-2000 maglev magnets.,
The question is: maglev is commercially used. True or False?
The answer is: false
it appears that the super-conducting maglev system is technically ready to be used commercially as a very high-speed, large-capacity transportation system.
The question is: maglev is commercially utilize. True or False?
The answer is true.
Table 5: This example illustrates the TextBugger attack instance. The portion surrounded by a gray background represents our test example, and the text in blue font indicates the words that were altered in the TextBugger attack question.
Unperturbed actual statistics about the deterrent value of capital punishment are not available because it is impossible to know who may have been deterred from committing a crime.
The question is: capital punishment is a deterrent to crime. True or False?
The answer is false.
capital punishment is a catalyst for more crime.
The question is: capital punishment is a deterrent to crime. True or False?
The answer is false.
rather than deterring crime, capital punishment actually increases the level of brutality in society.,
The question is: capital punishment is a deterrent to crime. True or False?
The answer is false.
, the death penalty is not a deterrent.
The question is: capital punishment is a deterrent to crime. True or False?
The answer is false.
a closely divided u.s. supreme court said on thursday its 2002 ruling that juries and not judges must impose a death sentence applies only to future cases, a decision that may affect more than 100 death row inmates.
The question is: the supreme court decided that only judgescan impose the death sentence. True or False?
The answer is false.
Perturbed actual statistics about the deterrent value of capital punishment are not available because it is impossible to know who may have been deterred from committing a crime.
The question is: capital punishment is a deterrent to crime. True or False?
The answer is false.
capital punishment is a catalyst for more crime.
The question is: capital punishment is a deterrent to crime. True or False?
The answer is false.
rather than deterring crime, capital punishment actually increases the level of brutality in society.
The question is: capital punishment is a deterrent to crime. True or False?
The answer is false.
, the death penalty is not a deterrent.
The question is: capital punishment is a deterrent to crime. True or False?
The answer is false.
a closely divided u.s. supreme court said on thursday its 2002 ruling that juries and not judges must impose a death sentence applies only to future cases, a decision that may affect more than 100 death row inmates.
The question is: the high court decision that only magistrates can impose the death condemnation. True or False?
The answer is true.
Table 6: This example illustrates the TextFooler attack instance. The portion surrounded by a gray background represents our test example, and the text in blue font indicates the words that were altered in the TextFooler attack question.
Unperturbed gold mining operations in california and nevada use cyanide to extract the precious metal.
The question is: cyanide is used in gold mining. True or False?
The answer is: true
access to the underground workings at the la camorra mine is via a ramp from the surface, excavated at a -15% grade and connecting numerous levels.
The question is: la camorra is a mine. True or False?
The answer is: true
the mine would operate nonstop seven days a week and use tons of cyanide each day to leach the gold from crushed ore.
The question is: a weak cyanide solution is poured over it to pull the gold from the rock. True or False?
The answer is: false
the dam covering what used to be its football pitch fill up with millions of tons of poisonous mining waste.
The question is: mine waste-water poses an environmental hazard. True or False?
The answer is: false
known as “heap leach” mining, the method has become popular in the last decade because it enables microscopic bits of gold to be economically extracted from low-grade ore.
The question is: the mining industry uses a method known as heap leaching. True or False?
The answer is true.
Perturbed gold mining operations in california and nevada use cyanide to extract the precious metal.
The question is: cyanide is used in gold mining. True or False?
The answer is: true
access to the underground workings at the la camorra mine is via a ramp from the surface, excavated at a -15% grade and connecting numerous levels.
The question is: la camorra is a mine. True or False?
The answer is: true
the mine would operate nonstop seven days a week and use tons of cyanide each day to leach the gold from crushed ore.
The question is: a weak cyanide solution is poured over it to pull the gold from the rock. True or False?
The answer is: false
the dam covering what used to be its football pitch fill up with millions of tons of poisonous mining waste.
The question is: mine waste-water poses an environmental hazard. True or False?
The answer is: false
known as “heap leach” mining, the method has become popular in the last decade because it enables microscopic bits of gold to be economically extracted from low-grade ore.
The question is: the mining industry uses a method known as heap leachmanufacturing. True or False?
The answer is false.
Table 7: This example illustrates the BERT-Attack attack instance. The portion surrounded by a gray background represents our test example, and the text in blue font indicates the words that were altered in the BERT-Attack attack question.
Unperturbed in other words, with its 2 million inhabitants, slovenia has only 5.5 thousand professional soldiers.
The question is: slovenia has 5.5 million inhabitants. True or False?
The answer is:false
nicholas burns stated in an exclusive interview for ”utrinski vesnik” from pristina that it would be disgraceful if athens puts a veto on macedonia’s application for membership in the european union and nato.
The question is: greece and macedonia are in dispute over name. True or False?
The answer is: false
greece objects to the neighbouring skopje government using the name macedonia, saying it implies claims on a greek province of the same name.
The question is: greece and macedonia are in dispute over name. True or False?
The answer is: true
in other words, with its 2 million inhabitants, slovenia has only 5.5 thousand professional soldiers.
The question is: slovenia has 2 million inhabitants. True or False?
The answer is: true
the croatian intent is even more problematic because the border between slovenian and croatian territorial waters has not yet been established. the dispute about this border began in 1991 when both countries became independent.
The question is: there is a territorial waters dispute. True or False?
The answer is true.
Perturbed in other words, with its 2 million inhabitants, slovenia has only 5.5 thousand professional soldiers.
The question is: slovenia has 5.5 billion inhabitants. True or False?
The answer is:false
nicholas burns stated in an exclusive interview for ”utrinski vesnik” from pristina that it would be disgraceful if athens puts a veto on macedonia’s application for membership in the european union and nato.
The question is: greece and macedonia are in dispute over name. True or False?
The answer is: false
greece objects to the neighbouring skopje government using the name macedonia, saying it implies claims on a greek province of the same name.
The question is: greece and macedonia are in dispute over name. True or False?
The answer is: true
in other words, with its 2 million inhabitants, slovenia has only 5.5 thousand professional soldiers.
The question is: slovenia has 2 million inhabitants. True or False?
The answer is: true
the croatian intent is even more problematic because the border between slovenian and croatian territorial waters has not yet been established. the dispute about this border began in 1991 when both countries became independent.
The question is: there is a territorial waters dispute. True or False?
The answer is false.
Table 8: This example illustrates the AdvICL attack instance. The portion surrounded by a gray background represents our test example, and the text in blue font indicates the words that were altered in the AdvICL attack question.
Unperturbed money raised from the sale will go into a trust for hepburn’s family.
The question is:proceeds go to hepburn’s family. True or False?
The answer is:true
treasures belonging to hollywood legend katharine hepburn have raised £3.2m at a two-day auction in america.
The question is: a two-day auction of property belonging to actress katharine hepburn brought in 3.2 million pounds. True or False?
The answer is: true
a two-day auction of property belonging to actress katharine hepburn brought in 3.2 million pounds.
The question is: a two-day auction of property belonging to actress katharine hepburn brought in £3.2m. True or False?
The answer is: true
hepburn, a four-time academy award winner, died last june in connecticut at age 96.
The question is: hepburn, who won four oscars, died last june aged 96. True or False?
The answer is: true
hepburn’s family will receive the proceeds from the sale.
The question is: proceeds go to hepburn’s family. True or False?
The answer is true.
Perturbed money raised from the sale will go into a trust for hepburn’s family.
The question is:proceeds go to hepburn’s family. True or False?
The answer is: false
treasures belonging to hollywood legend katharine hepburn have raised £3.2m at a two-day auction in america.
The question is: a two-day auction of property belonging to actress katharine hepburn brought in 3.2 million pounds. True or False?
The answer is: true
a two-day auction of property belonging to actress katharine hepburn brought in 3.2 million pounds.
The question is: a two-day auction of property belonging to actress katharine hepburn brought in £3.2m. True or False?
The answer is: false
hepburn, a four-time academy award winner, died last june in connecticut at age 96.
The question is: hepburn, who won four oscars, died last june aged 96. True or False?
The answer is: true
hepburn’s family will receive the proceeds from the sale.
The question is: proceeds go to hepburn’s family. True or False?
The answer is false.
Table 9: This example illustrates the Swap-Labels attack instance. The portion surrounded by a gray background represents our test example, and the text in blue font indicates the words that were altered in the Swap-Labels attack question.
Unperturbed herceptin was already approved to treat the sickest breast cancer patients, and the company said, monday, it will discuss with federal regulators the possibility of prescribing the drug for more breast cancer patients.
The question is: herceptin can be used to treat breast cancer. True or False?
The answer is:true
only 14 percent of u.s. mothers exclusively breast-feed their babies for the minimum recommended six months.
The question is: there are many benefits from breast-feeding. True or False?
The answer is: false
new research shows there has been a sharp increase in disfiguring skin cancers, particularly in women under the age of 40, providing more evidence that young people are not heeding warnings about the dangers of tanning.
The question is: tanning may cause skin cancers. True or False?
The answer is: true
weinstock painstakingly reviewed dozens of studies for evidence of any link between sunscreen use and either an increase or decrease in melanoma.
The question is: skin cancer numbers increase. True or False?
The answer is: false
a compound in breast milk has been found to destroy many skin warts, raising hopes it also might prove effective against cervical cancer and other lethal diseases caused by the same virus
The question is: breast milk may help fight cervical cancer. True or False?
The answer is true.
Perturbed herceptin was already approved to treat the sickest breast cancer patients, and the company said, monday, it will discuss with federal regulators the possibility of prescribing the drug for more breast cancer patients.
The question is: herceptin can be used to treat breast cancer. True or False?
The answer is: false
only 14 percent of u.s. mothers exclusively breast-feed their babies for the minimum recommended six months.
The question is: there are many benefits from breast-feeding. True or False?
The answer is: false
new research shows there has been a sharp increase in disfiguring skin cancers, particularly in women under the age of 40, providing more evidence that young people are not heeding warnings about the dangers of tanning.
The question is: tanning may cause skin cancers. True or False?
The answer is: true
weinstock painstakingly reviewed dozens of studies for evidence of any link between sunscreen use and either an increase or decrease in melanoma.
The question is: skin cancer numbers increase. True or False?
The answer is: true
a compound in breast milk has been found to destroy many skin warts, raising hopes it also might prove effective against cervical cancer and other lethal diseases caused by the same virus
The question is: breast milk may help fight cervical cancer. True or False?
The answer is false.
Table 10: This example illustrates the Swap-Labels (Fix) attack instance. The portion surrounded by a gray background represents our test example, and the text in blue font indicates the words that were altered in the Swap-Labels (Fix) attack question.
Unperturbed rabies virus infects the central nervous system, causing encephalopathy and ultimately death. early symptoms of rabies in humans are nonspecific, consisting of fever, headache, and general malaise.
The question is: rabies is fatal in humans. True or False?
The answer is:true
rabies is a viral disease of mammals and is transmitted primarily through bites. annually, 7,000 to 8,000 rabid animals are detected in the united states, with more than 90 percent of the cases in wild animals.
The question is: rabies is fatal in humans. True or False?
The answer is: false
former tennis star vitas gerulaitis died in such fashion in september, from a lethal carbon monoxide buildup related to the faulty installation of a propane heater.
The question is: vitas gerulaitis died of carbon monoxide poisoning. True or False?
The answer is: true
more than 6,400 migratory birds and other animals were killed in nevada by drinking water in the cyanide-laced ponds produced by gold mining operations.
The question is: animals have died by the thousands from drinking at cyanide-laced holding ponds. True or False?
The answer is: true
a farmer who was in contact with cows suffering from bse – the so-called mad cow disease – has died from what is regarded as the human form of the disease.
The question is: bovine spongiform encephalopathy is another name for the “mad cow disease”. True or False?
The answer is true.
Perturbed The further into summer we go, the more California’s winter swells seem like a distant memory. Do you remember? Tons of rain, a few massive days up and down the coast, and more rain. … The latter half’s roping barrels will make you want to throw on your booties, crank the heat, and grab your step-up.
The question is: A motorcyclist struck a cow and suffered fatal injuries in Pierce County on Saturday morning, authorities said. True or False?
The answer is: true
most of the mines are in arid areas and animals searching for water are attracted to the cyanide-laced holding ponds that are an integral part of the mining operations.
The question is: animals have died by the thousands from drinking at cyanide-laced holding ponds. True or False?
The answer is: false
the unconfirmed case in dundee concerns a rabies-like virus known only in bats.
The question is: a case of rabies was confirmed. True or False?
The answer is: false
this case of rabies in western newfoundland is the first case confirmed on the island since 1989.
The question is: a case of rabies was confirmed. True or False?
The answer is: true.
a farmer who was in contact with cows suffering from bse – the so-called mad cow disease – has died from what is regarded as the human form of the disease.
The question is: bovine spongiform encephalopathy is another name for the “mad cow disease”. True or False?
The answer is false.
Table 11: This example illustrates the Irrelevant attack instance. The portion surrounded by a gray background represents our test example, and all demonstrations have been replaced by irrelevant attacks.
Shots Method Clean Test Sample Attack Demonstration Attack Datastore Attack
TB TF BA AdvICL S-L S-L (Fix) Irr.
$m=2$ ICL 94.65±0.85 57.57±2.14 49.81±0.96 25.46±0.35 68.33±6.72 25.11±14.34 74.29±2.24 92.01±0.89
$k$NN-ICL 88.3±7.85 49.08±2.88 47.17±1.15 33.29±6.29 24.15±2.17 30.11±4.47 30.28±1.89 89.22±6.47
$\text{R}_{\text{BM25}}$-ICL 90.48 43.12 44.04 36.35 73.62 6.88 55.3 89.91
$\text{R}_{\text{SBERT}}$-ICL 92.43 56.42 51.61 43.46 74.33 7.43 59.20 89.45
$\text{R}_{\text{Instructor}}$-ICL 94.95 58.49 53.56 41.06 71.29 16.9 57.39 92.66
$m=4$ ICL 94.49±1.1 58.53±1.89 51.34±1.27 25.92±0.3 69.33± 16.11±9.49 61.29±3.24 90.48±0.64
$k$NN-ICL 90.25±5.47 52.03±4.04 50.3±1.95 34.48±2.99 25.28±4.18 30.11±4.47 34.28±4.90 91.01±2.12
$\text{R}_{\text{BM25}}$-ICL 93.0 53.67 49.66 42.32 76.83 1.49 57.8 91.86
$\text{R}_{\text{SBERT}}$-ICL 87.84 52.64 46.67 43.12 71.28 3.27 52.73 83.69
$\text{R}_{\text{Instructor}}$-ICL 92.78 59.86 53.56 42.32 72.09 4.77 51.84 89.89
$m=8$ ICL 92.66±3.18 58.11±5.01 52.18±2.29 29.28±1.56 61.49±5.23 7.11±8.49 34.29±8.24 91.09±3.6
$k$NN-ICL 92.01±2.13 54.28±2.57 50.19±1.26 35.28±2.38 19.28±4.28 30.11±4.47 31.28±3.92 92.01±1.53
$\text{R}_{\text{BM25}}$-ICL 94.61 53.56 51.15 40.37 79.82 0.11 34.17 93.81
$\text{R}_{\text{SBERT}}$-ICL 95.18 63.76 63.76 41.63 69.28 0.34 31.56 89.38
$\text{R}_{\text{Instructor}}$-ICL 94.84 58.49 63.90 39.22 68.08 0.44 29.10 88.46
$m=16$ ICL 92.12±1.06 56.69±5.07 49.39±4.7 28.29±3.46 60.67±5.2 4.11±4.27 18.30 78.02±21.45
$k$NN-ICL 93.65±1.49 53.97±1.79 52.37±1.74 36.57±1.28 15.79±6.22 30.11±4.47 31.99±3.20 92.54±2.56
$\text{R}_{\text{BM25}}$-ICL 94.61 56.77 50.25 38.76 70.5 0.33 15.14 91.72
$\text{R}_{\text{SBERT}}$-ICL 95.07 60.32 53.29 41.51 64.08 0.00 12.47 92.72
$\text{R}_{\text{Instructor}}$-ICL 94.38 59.06 53.18 41.28 62.08 0.00 11.99 92.61
Table 12: Complete results for the SST-2 dataset. The term Clean refers to the benign accuracy before attacks; attack accuracy is provided under each attack column.
Shots Method Clean Test Sample Attack Demonstration Attack Datastore Attack
TB TF BA AdvICL S-L S-L (Fix) Irr.
$m=2$ ICL 68.47±5.98 18.65±7.33 20.46±8.66 3.01±2.41 60.73±3.87 20.22±9.07 48.26±11.07 67.27±4.41
$k$NN-ICL 61.25±6.47 23.47±18.96 25.39±16.97 4.28±3.11 3.28±2.77 29.96±2.96 31.05±2.73 64.86±7.09
$\text{R}_{\text{BM25}}$-ICL 70.76 14.08 20.58 3.25 60.38 13.78 56.68 62.82
$\text{R}_{\text{SBERT}}$-ICL 70.04 11.91 21.30 1.81 59.39 11.49 36.29 59.57
$\text{R}_{\text{Instructor}}$-ICL 71.48 14.83 18.05 1.81 59.40 9.09 37.02 68.59
$m=4$ ICL 70.88±2.12 17.09±6.32 20.7±6.55 2.65±1.99 61.27±5.39 12.03±3.36 43.68±10.4 66.19±3.51
$k$NN-ICL 64.26±6.25 9.93±5.87 15.73±3.33 3.18±2.62 4.27±2.19 25.15±3.99 26.35±4.34 59.33±6.09
$\text{R}_{\text{BM25}}$-ICL 68.59 15.16 27.80 3.18 61.22 8.66 34.77 59.21
$\text{R}_{\text{SBERT}}$-ICL 71.84 14.44 16.97 0.72 58.27 9.20 31.29 60.29
$\text{R}_{\text{Instructor}}$-ICL 71.48 19.86 18.05 2.76 55.38 10.23 37.78 61.48
$m=8$ ICL 73.04±1.27 12.75±2.4 17.81±3.76 1.56±0.56 62.58±1.37 6.14±1.81 30.45±3.36 64.98±7.19
$k$NN-ICL 70.52±3.90 6.62±2.56 13.84±5.28 3.22±1.41 2.73±1.19 25.63±3.82 25.99±3.77 65.7±4.7
$\text{R}_{\text{BM25}}$-ICL 71.48 22.38 26.35 6.86 58.06 5.42 28.16 60.65
$\text{R}_{\text{SBERT}}$-ICL 72.20 11.91 12.27 1.36 52.35 3.25 25.63 61.01
$\text{R}_{\text{Instructor}}$-ICL 73.29 15.88 20.22 2.44 52.71 6.86 24.92 65.70
$m=16$ ICL 73.53±0.75 8.9±2.18 11.91±1.88 1.2±0.21 59.21±8.90 4.17±1.62 24.26±3.26 68.95±2.37
$k$NN-ICL 70.88±3.51 14.44±7.05 16.97±6.03 1.32±0.98 1.73±0.22 24.89±2.91 25.12±3.02 69.67±3.82
$\text{R}_{\text{BM25}}$-ICL 64.15 20.94 41.52 18.41 38.27 3.83 27.87 58.84
$\text{R}_{\text{SBERT}}$-ICL 73.98 15.78 26.71 4.33 54.29 4.87 26.80 64.26
$\text{R}_{\text{Instructor}}$-ICL 69.33 15.16 23.83 1.44 56.86 7.12 28.12 64.26
Table 13: Complete results for the RTE dataset. The term Clean refers to the benign accuracy before attacks; attack accuracy is provided under each attack column.
Shots Method Clean Test Sample Attack Demonstration Attack Datastore Attack
TB TF BA AdvICL S-L S-L (Fix) Irr.
$m=2$ ICL 91.97±0.21 57.4±1.65 49.67±0.32 24.67±0.32 75.33±3.72 23.11±12.36 48.57±0.64 86.93±7.05
$k$NN-ICL 90.9±0.42 55.95±0.92 49.67±1.86 33.67±4.11 23.77±1.17 25.38±8.47 26.88±7.47 87.12±12.01
$\text{R}_{\text{BM25}}$-ICL 89.6 44.38 47.92 26.83 73.26 16.1 75.3 88.9
$\text{R}_{\text{SBERT}}$-ICL 90.4 57.4 53.6 29.22 70.33 9.7 78.8 89.7
$\text{R}_{\text{Instructor}}$-ICL 91.3 42.9 32.6 28.45 71.43 8.4 86.9 89.8
$m=4$ ICL 92.07±0.32 58.0±2.6 50.87±1.21 28.97±0.5 73.29±4.02 14.11±8.36 45.37±2.66 92.13±0.8
$k$NN-ICL 92.3±1.04 56.93±1.5 49.67±1.86 31.67±3.20 21.33±4.18 24.76±3.40 26.12±3.51 91.22±4.01
$\text{R}_{\text{BM25}}$-ICL 92.0 48.2 49.8 29.01 71.38 3.3 61.0 90.3
$\text{R}_{\text{SBERT}}$-ICL 91.6 56.6 54.2 33.28 70.24 4.5 51.8 91.4
$\text{R}_{\text{Instructor}}$-ICL 91.6 49.5 40.7 31.47 69.11 2.8 78.4 91.7
$m=8$ ICL 92.67±0.35 58.93±2.63 49.37±1.07 30.1±0.5 71.49±6.23 5.11±4.28 22.77±14.96 92.53±0.49
$k$NN-ICL 91.77±0.32 55.43±3.52 48.47±2.08 31.25±1.20 19.28±3.21 25.87±2.43 24.87±5.43 89.22±3.01
$\text{R}_{\text{BM25}}$-ICL 92.5 57.4 50.1 31.5 69.28 0.3 33.4 91.1
$\text{R}_{\text{SBERT}}$-ICL 92.4 56.8 49.8 32.78 65.76 0.2 36.9 92.4
$\text{R}_{\text{Instructor}}$-ICL 92.6 57.9 48.8 31.28 68.28 1.8 32.1 92.2
$m=16$ ICL 92.77±0.45 57.8±0.53 50.53±0.59 29.57±0.5 69.67±3.1 4.11±2.27 21.77±6.96 93.0±0.4
$k$NN-ICL 92.13±0.29 55.63±1.86 48.9±1.35 32.25±0.34 17.79±2.12 22.87±3.61 22.87±2.43 87.22±2.01
$\text{R}_{\text{BM25}}$-ICL 93.5 59.6 51.0 29.01 65.5 4.4 20.34 92.3
$\text{R}_{\text{SBERT}}$-ICL 93.0 55.7 48.8 31.47 61.12 6.6 19.85 92.5
$\text{R}_{\text{Instructor}}$-ICL 93.1 57.2 49.3 33.28 63.33 5.3 16.38 92.5
Table 14: Complete results for the MR dataset. The term Clean refers to the benign accuracy before attacks; attack accuracy is provided under each attack column.
Shots Method Clean Test Sample Attack Demonstration Attack Datastore Attack
TB TF BA AdvICL S-L S-L (Fix) Irr.
$m=2$ ICL 93.35±0.96 66.04±2.17 47.61±2.27 27.48±2.47 76.21±4.27 16.04±13.28 63.03±4.52 88.12±1.26
$k$NN-ICL 71.87±11.2 29.4±2.38 29.4±2.38 29.4±2.38 27.15±6.27 15.28±2.38 17.90±1.09 64.4±6.38
$\text{R}_{\text{BM25}}$-ICL 92.29 63.03 51.06 0 69.62 24.20 80.59 91.29
$\text{R}_{\text{SBERT}}$-ICL 92.02 62.23 51.60 32.71 71.34 8.24 82.18 92.0
$\text{R}_{\text{Instructor}}$-ICL 93.09 63.56 58.51 39.36 72.29 1.86 90.69 92.1
$m=4$ ICL 93.26±0.85 61.79±7.05 45.92±2.47 28.28±0.67 71.33±2.01 1.68±2.47 47.87±1.88 89.2±1.44
$k$NN-ICL 73.87±8.40 29.4±2.38 29.4±2.38 29.4±2.38 26.28±3.18 17.28±1.29 19.4±4.29 67.4±1.38
$\text{R}_{\text{BM25}}$-ICL 91.49 65.65 45.48 0 67.24 1.86 61.70 90.34
$\text{R}_{\text{SBERT}}$-ICL 93.09 56.38 48.67 32.18 65.28 0.53 69.68 91.2
$\text{R}_{\text{Instructor}}$-ICL 93.27 62.50 60.37 42.82 64.09 0.00 88.03 90.47
$m=8$ ICL 91.31±3.31 58.86±7.22 43.44±4.25 26.33±1.88 70.49±3.25 0.09±0.16 20.21±16.51 90.2±0.96
$k$NN-ICL 91.87±4.29 29.4±2.38 29.4±2.38 29.4±2.38 20.19±2.11 18.28±2.66 18.9±2.37 69.4±1.22
$\text{R}_{\text{BM25}}$-ICL 93.09 65.96 48.94 31.91 62.18 0.00 40.69 86.29
$\text{R}_{\text{SBERT}}$-ICL 93.62 64.36 50.27 26.60 61.97 0.00 35.28 89.16
$\text{R}_{\text{Instructor}}$-ICL 93.88 61.44 56.91 43.35 60.26 0.00 32.97 90.11
$m=16$ ICL 84.31±12.73 53.37±10.37 41.58±5.53 28.9±2.03 66.67±2.5 3.1±0.15 61.88±21.06 84.1±0.96
$k$NN-ICL 83.29±2.1 29.4±2.38 29.4±2.38 29.4±2.38 18.79±1.99 19.28±1.83 20.2±1.97 74.4±0.27
$\text{R}_{\text{BM25}}$-ICL 93.09 68.62 60.11 38.30 59.34 9.31 85.90 87.28
$\text{R}_{\text{SBERT}}$-ICL 93.88 69.95 54.52 30.05 61.27 6.12 74.47 90.32
$\text{R}_{\text{Instructor}}$-ICL 94.00 64.10 60.90 45.21 63.11 4.21 77.13 87.10
Table 15: Complete results for the CR dataset. The term Clean refers to the benign accuracy before attacks; attack accuracy is provided under each attack column.
Shots Method Clean Test Sample Attack Demonstration Attack Datastore Attack
TB TF BA AdvICL S-L S-L (Fix) Irr.
$m=2$ ICL 51.07±4.14 19.07±4.92 21.83±5.79 24.22±4.79 42.10±6.54 10.43±0.32 35.10±4.72 50.4±3.38
$k$NN-ICL 53.07±5.28 17.03±5.29 22.03±2.26 21.38±5.22 39.14±3.98 19.03±8.22 28.48±3.18 49.07±4.67
$\text{R}_{\text{BM25}}$-ICL 58.4 13.6 16.5 30.18 28.4 2.8 28.67 49.6
$\text{R}_{\text{SBERT}}$-ICL 55.7 12.6 20.8 26.43 30.42 0.8 26.1 53.7
$\text{R}_{\text{Instructor}}$-ICL 53.2 14.6 20.4 25.71 31.74 1.6 22.5 54.1
$m=4$ ICL 54.17±2.19 17.33±2.61 22.47±2.0 27.68±3.25 38.23±4.66 8.03±0.06 32.10±6.15 52.7±1.59
$k$NN-ICL 55.07±2.72 16.12±3.87 25.13±3.26 23.4±3.19 30.14±1.57 22.03±6.31 26.7±3.75 51.46±3.67
$\text{R}_{\text{BM25}}$-ICL 54.8 17.4 19.9 29.75 34.57 1.1 24.36 55.4
$\text{R}_{\text{SBERT}}$-ICL 57.5 13.0 16.8 28.84 33.81 0.5 24.08 55.8
$\text{R}_{\text{Instructor}}$-ICL 54.3 16.6 14.2 27.79 32.68 0.3 20.47 56.5
$m=8$ ICL 53.63±2.99 16.9±4.33 22.57±4.94 31.72±2.17 36.21±3.27 6.0±0.0 30.12±1.48 52.7±1.75
$k$NN-ICL 55.89±3.61 19.22±2.62 26.77±2.52 27.22±3.01 27.22±1.25 21.02±4.18 24.59±3.11 50.22±1.90
$\text{R}_{\text{BM25}}$-ICL 57.3 18.1 23.4 33.82 35.27 2.3 20.69 57.1
$\text{R}_{\text{SBERT}}$-ICL 57.8 17.9 25.97 32.79 36.41 0.4 21.07 56.2
$\text{R}_{\text{Instructor}}$-ICL 54.3 18.2 26.1 34.72 33.76 0.0 16.47 52.9
$m=16$ ICL 53.3±3.13 17.73±4.9 22.63±2.87 35.72±1.88 38.32±5.81 4.3±2.0 29.48±2.46 51.8±0.82
$k$NN-ICL 57.89±1.34 20.79±1.93 25.68±1.94 36.17±1.05 18.22±6.31 23.48±1.23 25.12±4.15 49.37±0.63
$\text{R}_{\text{BM25}}$-ICL 59.1 20.2 21.78 38.09 38.02 1.1 19.72 52.3
$\text{R}_{\text{SBERT}}$-ICL 58.9 21.5 23.4 36.4 37.45 0.2 20.78 54.3
$\text{R}_{\text{Instructor}}$-ICL 55.5 22.4 26.9 33.21 33.47 0.4 18.93 53.8
Table 16: Complete results for the MNLI dataset. The term Clean refers to the benign accuracy before attacks; attack accuracy is provided under each attack column.
Shots Method Clean Test Sample Attack Demonstration Attack Datastore Attack
TB TF BA AdvICL S-L S-L (Fix) Irr.
$m=2$ ICL 69.87±9.93 19.6±10.05 27.53±7.76 25.13±5.77 60.73±4.87 0.13±0.23 0.07±0.12 59.12±4.93
$k$NN-ICL 71.87±11.2 18.4±7.38 26.4±8.92 31.48±5.89 33.28±6.77 30.28±2.38 31.90±4.09 64.4±6.38
$\text{R}_{\text{BM25}}$-ICL 70.2 26 10.6 14.4 60.28 0.2 30.6 65.3
$\text{R}_{\text{SBERT}}$-ICL 84.0 35.6 11.6 41.8 61.34 2.0 41.0 78.12
$\text{R}_{\text{Instructor}}$-ICL 75.6 30.8 11.8 39.0 62.44 1.8 46.0 72.10
$m=4$ ICL 70.53±9.56 22.67±8.62 26.27±11.44 24.67±8.77 59.27±5.39 21.87±2.53 12.53±9.1 65.22±3.78
$k$NN-ICL 73.87±8.40 27.4±4.38 31.2±5.22 36.1±4.29 24.27±5.2 31.98±1.29 34.41±3.29 67.4±1.38
$\text{R}_{\text{BM25}}$-ICL 82.8 5.0 13.6 20.2 55.32 27 52.2 79.2
$\text{R}_{\text{SBERT}}$-ICL 84.8 43.2 18.4 44.0 54.76 28.2 54.6 81.4
$\text{R}_{\text{Instructor}}$-ICL 82.0 34.6 22.0 38.4 51.29 30.4 58.4 80.6
$m=8$ ICL 76.07±5.88 29.6±6.55 32.33±6.93 29.93±6.64 61.85±3.17 26.2±3.68 45.2±6.32 69.84±2.11
$k$NN-ICL 77.87±4.29 31.4±5.22 35.4±3.92 38.92±2.76 12.73±2.18 32.13±3.66 36.9±2.37 69.4±1.22
$\text{R}_{\text{BM25}}$-ICL 79.2 39.6 39.0 34.0 56.3 35.2 71.6 76.2
$\text{R}_{\text{SBERT}}$-ICL 87.8 49.4 47.8 46.4 57.41 41.2 79.0 79.8
$\text{R}_{\text{Instructor}}$-ICL 86.5 47.2 25.2 45.4 59.42 47.6 83.6 80.2
$m=16$ ICL 83.53±3.41 38.07±3.23 42.2±5.03 39.53±3.82 59.21±8.90 32.2±1.78 65.2±3.32 71.21±1.33
$k$NN-ICL 83.29±2.1 39.4±3.28 41.39±3.21 42.1±3.12 10.73±2.23 33.28±4.83 36.2±1.97 74.4±0.27
$\text{R}_{\text{BM25}}$-ICL 82.0 42.4 40.4 37.0 58.27 41.2 77.2 80.2
$\text{R}_{\text{SBERT}}$-ICL 90.0 54.8 56.8 48.4 56.13 43.8 84.2 87.4
$\text{R}_{\text{Instructor}}$-ICL 89.6 52.0 35.0 47.4 56.68 42.4 86.6 88.2
Table 17: Complete results for the TREC dataset. The term Clean refers to the benign accuracy before attacks; attack accuracy is provided under each attack column.
Defence Shots Clean Accuracy TB TF BA
$\hookrightarrow$ No Defence 8 71.48 22.38 26.35 6.86
Augmentation
$\hookrightarrow$ Random Addition1 8 72.01 23.22 23.21 7.29
$\hookrightarrow$ Random Deletion1 8 69.18 23.71 25.75 9.71
DARD
$\hookrightarrow$ R-ICL (BM25) 8 74.39 31.02 36.13 15.38
$\hookrightarrow$ R-ICL (SBERT) 8 74.16 30.53 41.19 20.32
$\hookrightarrow$ R-ICL (Instructor) 8 71.38 22.13 27.26 11.20
Adversarial Training 8 77.22 29.17 44.82 17.59
Table 18: DARD for adversarial defences. We show the Clean Accuracy and the Attack Accuracy.

### v3

@@
-Evaluating the Adversarial Robustness of Retrieval-Based In-Context Learning for Large Language Models
+Evaluating and Safeguarding the Adversarial Robustness of Retrieval-Based In-Context Learning
 Abstract
-With the emergence of large language models, such as LLaMA and OpenAI GPT-3, In-Context Learning (ICL) gained significant attention due to its effectiveness and efficiency.
+With the emergence of large language models, such as LLaMA and GPT, In-Context Learning (ICL) gained significant attention due to its effectiveness and efficiency.
 However, ICL is very sensitive to the choice, order, and verbaliser used to encode the demonstrations in the prompt.
@@
 Large Language models (LLMs) are revolutionising the field of NLP (Brown et al., 2020; Wei et al., 2022; Touvron et al., 2023).
-However, this also raises concerns about their robustness and trustworthiness.
-Although significant efforts are put into aligning LLMs for safety purposes, recent works still show LLMs can be vulnerable to adversarial inputs and jailbreaking (Zou et al., 2023; Xu et al., 2023b; Zeng et al., 2024).
+However, this also raises concerns about their robustness (Xu et al., 2023c).
+Although significant efforts are put into aligning LLMs for safety purposes, recent works still show LLMs can be vulnerable to adversarial inputs and jailbreaking (Zou et al., 2023; Xu et al., 2023b; Zeng et al., 2024; Longpre et al., 2024).
 As a result, it is essential to understand LLM security properties to guide the direction of LLMs that are secure and robust.
@@
 Prior work—which we review in Section 2—shows that LLMs can still be vulnerable to different types of input perturbations, such as task instructions (Zhu et al., 2023; Sun et al., 2023) or inputs (Wang et al., 2023a); however, the robustness of ICL-based methods to such perturbations is still under-explored.
-We address (in Section 4) the following question: How sensitive are vanilla and retrieval-based ICL methods to perturbations to their test samples or demonstration instances, which is particularly important given that the primary motivation of ICL is to facilitate few-shot learning without training; should any methods exhibit excessive sensitivity, this would limit their applicability.
+We address (cf.
+Section 4) the following question: How sensitive are vanilla and retrieval-based ICL methods to perturbations to their test samples or demonstration instances, which is particularly important given that the primary motivation of ICL is to facilitate few-shot learning without training; should any methods exhibit excessive sensitivity, this would limit their applicability.
 To investigate this problem, we employ commonly used perturbation attacks—namely character-level, word-level, and sentence-level attacks.
@@
 This encourages further investigation on enhancing model robustness.
-We further explore the advantages of utilising retrieval-based models to improve robustness (in Section 5), considering that existing defence strategies often rely on an adversarial training process (Li et al., 2021; Si et al., 2021).
+We further explore the advantages of utilising retrieval-based models to improve robustness (cf.
+Section 5), considering that existing defence strategies often rely on an adversarial training process (Li et al., 2021; Si et al., 2021).
 However, this process is memory intensive in the context of LLMs.
@@
 Therefore, we propose DARD, a training-free adversarial defence method which leverages adversarially perturbed samples to augment the retrieval pools, potentially improving the models’ robustness.
-The paper is summarized in the Figure 1.
+Figure 1 presents an overview of the paper.
 Our main contributions are summarized as follows.
@@
 It is important to note, however, that these experiments were mainly conducted using relatively small LMs such as BERT, RoBERTa.
-Recently, several works demonstrated LLM can still be vulnerable to these type of attacks, by showing the vulnerability of attacking the task instructions (PromptBench (Zhu et al., 2023)), or perform adversarial attacks on zero-shot settings and investigate the transferability of attacks from open-sourced models (i.e.
+Recently, several works demonstrated LLM can still be vulnerable to these type of attacks, by showing the vulnerability of attacking the task instructions (PromptBench (Zhu et al., 2023)), or performing adversarial attacks on zero-shot settings and investigating the transferability of attacks from open-sourced models (i.e.
 Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023)) to close-sourced (i.e.
@@
 Retrieval-based LLM
-Retrieval-based LLM has shown great success in lifting LLM performance and faithfulness and reducing hallucination, where the idea is to gather related articles or demonstrations from a collection of data.
+Retrieval-based LLM has shown great success in lifting LLM performance and faithfulness and reducing hallucination (Pan et al., 2023), where the idea is to gather related articles or demonstrations from a collection of data (Huang et al., 2023).
 Specifically, we are interested in whether they can be used to improve the robustness against perturbations (Lin & Lee, 2024).
@@
 (2022) — we replace 50% of the original demonstrations in the training data with Out-of-Distrubtion (OoD) examples.
-Further details and analyses are available in Appendix A.3.
+Further details and analyses about attack methods are available in Appendix A.4.
 4 Experiments
@@
 We include quantitative results in the Appendix B.2.
-4.3 Ablation Study
+4.3 Attack on Model Variants
 Figure 3: Attack Success Rate (%) for adversarial attacks across various models, based on experiments conducted on the RTE dataset with 8-shot demonstrations.
-The results are based on R-ICL and involve the aggregation of results from retrievers: BM25, SBERT, and Instructor.
+The results are based on R-ICL and the mean attack success rate among the three retrievers we used: BM25, SBERT, and Instructor.
 In this section, we extend our experiments to four other models: LLaMA-2-13B, larger version of 7B model; Vicuna-7B (Chiang et al., 2023), instruction-tuned from LLaMA-2-7B; Mistral-7B (Jiang et al., 2023) and Gemma-7B (Gemma Team et al., 2024), aiming to provide a view for models from different families to prevent potential bias posed within the pretraining data for LLaMA models.
@@
 Furthermore, we also demonstrate the potential of retrieval-based models in data augmentation strategies for adversarial defense.
-There are many open questions for future work.One of the most profound findings is that our study on attack transferability demonstrates that larger models generally exhibit better robustness; however, this does not necessarily hold true for Mixture of Experts (MoE) models.
+There are many open questions for future work.
+One of the most profound findings is that our study on attack transferability demonstrates that larger models generally exhibit better robustness; however, this does not necessarily hold true for Mixture of Experts (MoE) models.
 Given the recent surge in their popularity, it is crucial to investigate the conditions under which MoE models may fail and to understand the reasons behind their vulnerabilities.
+7 Limitations
+We view our work as an early exploration of the adversarial robustness of Retrieval-based in-context learning.
+Here, we only show how different types of attack can be applied to, but not why.
+Given the standard use of retrieval and in-context learning as inference methods, it is crucial to understand its deficiency.
+Future work should focus on certified robustness (Zhao et al., 2022; Xiang et al., 2024), which analyzes with theoretical proof the extent to which the model is unaffected by adversarial attacks.
+Acknowledgement
+Computations described in this research were supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh.
+We thank Giwon Hong and Yu Zhao for the insightful discussion at the beginning of this research work.
 Appendix A Experiment Settings
@@
 We conduct experiments on its released large-size model 333hkunlp/instructor-large.
-A.2 Templates and verbaliser for datasets
+A.2 The statistics of datasets
+Datasets No. of Labels Train/Test Size Avg. Length
+SST-2 2 67348/872 23.07
+RTE 2 2489/277 96.03
+MR 2 8530/1066 37.76
+CR 2 3394/376 30.80
+MNLI-mm 3 8832/982 55.90
+TREC 6 5452/500 19.59
+Table 3: Datasets Information
+The table 3 shows the number of categories for each classification dataset, the number of examples in the training and test sets, and the average sentence length.
+A.3 Templates and verbaliser for datasets
 See Table 4 for the used templates and verbaliser for mapping discrete labels into label space (Adopted from Min et al. (2022) and Lu et al. (2022)).
@@
 Table 4: Templates and verbaliser used across the experiments. These are minimum cases with only one demonstration example for illustration.
-A.3 Attack Details
+A.4 Attack Details
 We implemented our proposed attack methods and conducted all our experiments using the TextAttack framework (Morris et al., 2020), which is model-agnostic and can be easily tested on any model.
@@
   2. 2.
-    TextFooler: We also evaluated another popular word-level method, TextFooler (Jin et al., 2020), which is similar principle to TextBugger.
+    TextFooler: We also evaluated another popular word-level method, TextFooler (Jin et al., 2019), which is similar principle to TextBugger.
 However, TextFooler greedily searches for a large number of nearby words in the embedding space for each word, as long as they meet some constraints on embedding similarity and sentence quality.
@@
 The contamination rate is 50%.
-Examples of each attack are demonstrated in Table 5-11. Code will be released to facilitate further research in this field.
-A.4 Reproducability
+Examples of each attack are demonstrated in Table 6-12. Code will be released to facilitate further research in this field.
+A.5 Reproducability
 We conducted all attack experiments using either two A100-40GB GPUs or one A100-80GB GPU.
@@
 B.1 Main results
-We include the complete results breakdown in Table 12-17. The results shown in those tables are the Benign accuracy and Attack accuracy.
+We include the complete results breakdown in Table 13-18. The results shown in those tables are the Benign accuracy and Attack accuracy.
 B.2 Quantitative Results